### v2

# The behavioral spillover effect: Modeling behavioral interdependencies in multi-pathogen dynamics

Abstract
During the recent pandemic, a rise in COVID-19 cases was followed by a decline in influenza. In the absence of cross-immunity, a potential explanation for the observed pattern is behavioral: non-pharmaceutical interventions (NPIs) designed and promoted for one disease also reduce the spread of others. We study short-term and long-term dynamics of two pathogens where NPIs targeting one pathogen indirectly influence the spread of another – a phenomenon we term behavioral spillover. We examine how perceived risk of and response to one disease substantially alters the spread of other pathogens, revealing how waves of different pathogens emerge over time as a result of behavioral interdependencies and human response. Our analysis identifies the parameter space where two diseases simultaneously co-exist, and where shifts in prevalence occur. Our findings are consistent with observations from the COVID-19 pandemic, where NPIs contributed to significant declines in infections such as influenza, pneumonia, and Lyme disease.
Keywords: epidemic models, risk response, equilibria analysis, identifiability
1 Introduction
The epidemic modeling literature has evolved over the past century, incorporating various mathematical techniques to capture the complex dynamics of infectious disease spread within social and environmental contexts [32, 58, 4]. From early compartmental models based on ordinary differential equations to more sophisticated stochastic or network-based approaches, epidemiological models continuously advance to better represent real-world disease transmission patterns. The primary objective of these models are to enhance our understanding of disease dynamics [7], assess potential impacts of interventions to inform evidence-based policy-making [43, 66] and to offer short- and long-term projections [12, 57]. Recent advances include incorporation of change in human behavior and risk perception as well as mobility patterns in epidemiological models to further improve model-based insights [38].
Most studies focus on modeling the spread of a single virus, such as influenza or coronavirus [11, 39, 25]; however, multiple pathogens often co-circulate, with the disease with the higher prevalence shifting over time [48]. For example, influenza and respiratory syncytial virus frequently overlap during seasonal epidemics, with higher prevalence observed in colder months [41]. In contrast, Lyme disease is more common during warmer seasons when outdoor activities increase [46], and rhinoviruses peak in the spring and fall [48]. This raises an important question: to what extent are the dynamics of multiple co-circulating pathogens independent of each other? This question became especially important during early months of the COVID-19 pandemic when a large fraction of public health capacity was devoted to treating COVID-19 patients [59].
Despite the early concerns, an intriguing pattern emerged. As the coronavirus persisted over a relatively long period, a notable drop in influenza A and B cases was observed during the first two years of the pandemic [30, 62, 54, 33]. Interestingly, as the number of COVID-19 cases declined in late 2022, thanks to natural and vaccine-induced immunity, the number of influenza type-A cases skyrocketed (Figure 1). A potential explanation for the observed decline in influenza cases is behavioral: the non-pharmaceutical interventions (NPIs) implemented to control COVID-19, such as mask-wearing, social distancing, and reduced mobility, also limited the spread of other pathogens [62]. Since influenza transmission is partly aerosol-based, mask-wearing and other NPIs directly reduced influenza transmission as well. A second-order effect produced more complicated dynamics, related to the fact that during the two-year period of fewer influenza cases, little natural immunity was built against the virus, which possibly lead to a large wave of the disease once COVID-19 reached its endemic state.
This and other similar scenarios underscore the importance of analyzing interactions among multiple pathogens rather than relying solely on single-pathogen models. Single-pathogen approaches may overlook critical insights into competition, coexistence, and broader public health implications. Understanding the dynamics of multiple pathogens—including factors like cross-immunity and evolutionary processes—is crucial for capturing the complexity of real-world epidemics [71, 6]. For example, some studies have explored how partial immunity from one infection can influence the spread of other pathogens (e.g., [6, 24, 36]). Our focus, however, is distinct: we analyze how human behavior mediates the inter-dependencies among pathogens.
1.1 Cross immunity
Cross-immunity between different strains of the same pathogen or across different pathogen, referred to as heterologous immunity, has long been a topic of interest in immunology [71, 70]. Pathogens that trigger an immune response after infection can potentially provide at least partial cross-protection against strains of the same or closely related pathogens. This cross-protection can influence the spread of disease, with an epidemic of one strain suppressing the transmission of another [71]. Early in the COVID-19 pandemic, it was hypothesized that some level of cross-immunity might exist among coronaviruses, potentially reducing the spread or severity of SARS-CoV-2 infection [72].
The phenomenon of cross-immunity and evolution of various strains and mutations have been explored through mathematical modeling, e.g. [6, 24, 35, 18]. Bhattacharyya et al. [6] examine cross-immunity among respiratory viruses, focusing specifically on RSV, HPIV, and hMPV. Using compartmental models, they analyze patterns of laboratory-confirmed infections from the western United States between 2002 and 2014, arguing that closely related paramyxoviruses can suppress each other’s transmission through immune responses triggered by prior infections. In foundational work, Gog and Grenfell [24], showed how cross-immunity among multiple pathogens can produce a range of complex, yet realistic behavior.
It is notable that model-based studies of cross-immunity often modify conventional compartmental models to represent complicated states of immunity or partial immunity to one strain but susceptibility to others [36, 50]. However, based on recent pandemic experience, it is argued that such models often oversimplify or omit human behavior – either by assuming no variation in contact rates over the course of an epidemic or by formulating them exogenously [38]. In most contexts, fear of infection can lead to protective measures that suppress disease transmission, and as such epidemic models should better incorporate human behavior change during a course of a pandemic [22, 21, 68]. For example, during the recent COVID-19 pandemic, as case numbers rose, public risk perception also increased, prompting individuals to adjust their daily routines to reduce the risk of infection. It also influenced government policies, leading to more stringent measures [40]. Altogether, these responses influenced behavior with actions such as quarantining, social distancing, and mask-wearing. In turn, this response reduced transmission, leading to fewer cases, which subsequently lowered risk perception and contributed to a resurgence of the disease. This phenomenon illustrates a negative (balancing) feedback loop—one of several mechanisms that can lead to infection waves [56].
1.2 Modeling Human behavior
There are various approaches to modeling the spread of infectious disease, ranging from mechanistic models—including deterministic and stochastic compartmental models [7, 3]—to curve-fitting approaches such as machine learning [34, 14]. A comprehensive study of models used during the COVID-19 pandemic by Rahmandad et al. [57] finds that mechanistic compartmental models outperform curve-fitting approaches, particularly in long-term projections. Importantly, their results show that incorporating behavioral feedback—where contact rates adjust in response to changes in prevalence or mortality—substantially improves projection accuracy.
To capture this effect, Rahmandad et al. [57] propose the SEIRb model, which extends the standard Susceptible-Exposed-Infectious-Recovered (SEIR) framework by including a risk-responsive behavioral feedback mechanism. This allows infectivity to vary dynamically as human behavior adapts to the evolving state of the disease. As a result, even relatively simple mechanistic models can outperform more complex alternatives in long-term forecasting. In a related study, LeJeune et al. [37] develop a similar framework in which infectivity is explicitly modeled as a function of disease prevalence, with a delay reflecting lagged behavioral response. Their results demonstrate how such formulations improve the ability of models to capture realistic epidemic waves across data from 50 U.S. states and Washington, D.C.
These studies illustrate the growing emphasis on incorporating human behavior into epidemic models. More broadly, LeJeune et al. [38] classify such approaches into exogenous and endogenous formulations. Exogenous approaches treat behavioral changes as externally imposed parameter shifts, whereas endogenous approaches model quantities such as contact rate or infectivity as functions of the disease state itself. Within endogenous frameworks, several extensions have been proposed. These include modeling heterogeneous compliance by dividing the susceptible population into behavioral subgroups [1, 53], as well as representing individuals as payoff maximizers who balance economic costs against infection risks [60, 20]. Additional approaches incorporate the spread of fear, evolving social norms, and adherence fatigue [19, 55, 52], or use game-theoretic formulations to capture strategic behavioral adaptation [27, 9].
While these works highlight recent advancements in the consideration of cross-immunity and human behavior in epidemiological models, we emphasize here a critical gap which persists in the literature. On one hand, models of multi-pathogen dynamics often lack behavioral mechanisms; on the other hand, many behavioral epidemic models focus on single-pathogen contexts, overlooking how individual and collective responses influence outbreak dynamics. Particularly, compliance with NPIs can simultaneously influence the spread of multiple pathogens with similar transmission mechanisms.
1.3 Appearance in data
In real-world settings, public behavior is rarely pathogen-specific. For instance, increasing mask-wearing or social distancing in response to one disease can inadvertently suppress the transmission of others. This was clearly observed during the COVID-19 pandemic, when widespread NPI substantially reduced flu cases [30, 62, 54]. As shown in Figure 1, the United States has historically experienced annual flu seasons, typically peaking in late fall or early winter, until the onset of the recent COVID-19 pandemic. During the surge in COVID-19 cases in 2020 and 2021, influenza cases dramatically declined, with very few cases of Influenza A or B. Interestingly, as COVID-19 was brought under control and transitioned into an endemic phase, seasonal flu patterns re-emerged. In fact, Flu A cases surged during the 2022–2023 season with a peak that was not only substantially larger but also arrived earlier than those observed in pre-pandemic years. This pattern is often attributed to behavioral factors: compliance with NPIs not only curbed COVID-19 transmission but also suppressed flu outbreaks. Once such measures were relaxed, flu incidence rebounded. This heightened rebound can be partly attributed to a decline in population immunity, as the prolonged suppression of influenza circulation during the pandemic left a larger-than-usual proportion of the population with reduced or waning protection against circulating strains [10].
Both COVID-19 and influenza have similar transmission mechanisms which can be targeted by comparable prevention measures, likely contributing to the reduction in influenza cases observed during the initial outbreak of COVID-19. However, examining the literature and data of other infectious diseases with distinct transmission routes, such as COVID-19 and Lyme disease, we see that changes in behavior in response to COVID-19 may have also caused changes in the spread of non-respiratory diseases. For example, the lifestyle disruption that COVID-19 caused across the world resulted in changes in time spent outdoors, affecting potential for human exposure to tick-borne diseases [borșan2021recreational, 29, 44]. Some regions reported an increase in tick-borne encephalitis, possibly due to more time spent out-of-doors [31]. Other regions, though, saw a reduction in reported cases of Lyme disease, potentially resulting from disruptions in the healthcare workflow and healthcare-seeking habits which led to under-reporting of cases [29, 44, 51]. In either case, these correlations further support the hypothesis that disruptions in habits and routines - i.e., changes in public behavior - in response to one disease can affect the spread of another disease with an independent transmission pathway.
Figure 1: US Confirmed weekly cases of COVID-19 (blue, left y-axis) and Influenza A and B (red, right y-axis). Source: based on CDC Influenza data [8] and Johns Hopkins Dashboard COVID-19 data [16].
1.4 Beyond specific diseases
Similar dynamics are likely to arise in a variety of contexts and among pathogens with comparable transmission characteristics. For example, when individuals increase hand washing in response to a surge in one particular disease, they may also inadvertently reduce the transmission of other pathogens that spread through similar pathways. As such, modeling frameworks that incorporate behavioral coupling between co-circulating pathogens can offer unique insights into epidemic dynamics. These models can help answer the following important questions: How do general patterns of disease spread emerge when multiple pathogens are active and NPIs targeting one also reduce the transmission of another? How might these patterns change if such spillover effects are imperfect or only partially suppress the spread of a second disease? Can we identify ranges of reproductive numbers where one pathogen exceeds the other or where multiple pathogens can coexist? Addressing these questions is essential for improving our understanding of—and response to—the complex interplay between human behavior and the concurrent spread of multiple infectious diseases.
In this paper, we study the short-term and long-term dynamics of two pathogens spreading simultaneously, where NPIs targeting one pathogen can indirectly influence the spread of others—a phenomenon we term behavioral spillover. Generally, real-world infectious disease dynamics are influenced by many time-varying factors, including seasonality, policy changes, behavioral responses, emerging variants, environmental factors, and vaccination campaigns. The objective of this study, however, is not to reproduce the full complexity of epidemic dynamics, but rather a theoretical investigation: to isolate and analyze a specific mechanism—behaviorally mediated interactions between diseases—within a simple and analytically tractable framework. To that end, we present a mathematical model of two interacting pathogens and analyze their outbreak dynamics. The interaction is solely via human behavior; that is, no cross-immunity is considered, but we model how human risk response to one can fully or partially affect the spread of the other pathogen. We postulate that there are two regimes – persistence by only one pathogen or coexistence of both – and examine tipping points for change in patterns. This equilibrium is also influenced by the basic reproduction numbers of the pathogens. Over time, we observe a likelihood of periodic shift in which disease has higher prevalence between pathogens.
2 Materials and Methods
We develop a modified Susceptible-Infectious-Removed (SIR) model of the spread of two pathogens (denoted as $\{A,B\}$) with no cross-immunity and include endogenous behavioral responses as shown in Figure 2. To begin, we formulate the spread of each pathogen independently (excluding the green curved dashed lines) and then formulate the behavioral interconnection between the spread of the pathogens, considering both perfect and imperfect interconnection.
The core of our model structure is the simple SIR model with two modifications that include behavioral response and waning immunity. We refer to this family of models as SIRSb models to indicate the return of those in the recovered compartment to the susceptible compartment, and the inclusion of behavioral feedback that affects infectivity [37].
The system of ordinary differential equations, System (LABEL:eqn:SIRS), represents an SIRS model of disease $i$ (here we focus on two diseases, $i\in\{A,B\})$ with waning immunity given by
$\frac{dS_{i}}{dt}=-\beta_{i}(\cdot)S_{i}I_{i}+\frac{R_{i}}{\tau_{R}},$ (1)
$\frac{dI_{i}}{dt}=\beta_{i}(\cdot)S_{i}I_{i}-\frac{I_{i}}{\tau_{I}},$
$\frac{dR_{i}}{dt}=\frac{I_{i}}{\tau_{I}}-\frac{R_{i}}{\tau_{R}},$
$\frac{d\widetilde{I}_{i}}{dt}=\frac{I_{i}-\widetilde{I}_{i}}{\tau_{P}}$
$S_{i}+I_{i}+R_{i}=1,$
where $S_{i}$, $I_{i}$, and $R_{i}$ represent the proportion of individuals in the total population who are susceptible to disease $i$, infectious with disease $i$, and recovered from disease $i$, respectively. Notice that there is overlap among the two diseases in the population. The susceptible, infected, and recovered, populations with disease $A$ respectively, encompass the entire population, as do the susceptibles, infected, and recovered individuals with disease $B$. This stems from the assumption that all individuals have the potential to be infected with disease $A$, disease $B$, or both diseases simultaneously; in other words, individuals can become infected with both diseases without any biological cross-immunity. For example, a disease $A$ infectious individual is also simultaneously one of susceptible, infectious or recovered from disease $B$. Furthermore, infectious individuals do not change their behavior upon infection (e.g., no quarantine or isolation is included).
Susceptible individuals become infected with disease $i$ through interactions with infectious individuals of disease $i$ with transmission rate $\beta_{i}(\cdot)$, which can depend on other variables. Infectious individuals recover after an average period of $\tau_{I}$ days. Once recovered, individuals lose immunity and are once again susceptible after on average $\tau_{R}$ days.
Here, $\widetilde{I}_{i}$ represents the perceived number of disease $i$ infectious individuals, which follows the true value $I_{i}$ with an average delay $\tau_{P}$. Specifically, $\widetilde{I}_{i}$ adjusts toward $I_{i}$ at rate $1/\tau_{P}$, so that discrepancies between perceived and actual infections decay over time [47, 64]. This formulation corresponds to a standard exponential smoothing (information delay) process previously used in the literature to represent information flow (e.g., [65, 40]).
$S_{A}$ $I_{A}$ $R_{A}$ $\tilde{I}_{A}$ $\beta_{A}$$\dfrac{1}{\tau_{I}}$$\dfrac{1}{\tau_{R}}$$m_{A}$ $S_{B}$ $I_{B}$ $R_{B}$ $\tilde{I}_{B}$ $\beta_{B}$$\dfrac{1}{\tau_{I}}$$\dfrac{1}{\tau_{R}}$$m_{B}$$m_{A}$$m_{B}$
Figure 2: SIRS model with waning immunity and behavioral response for disease $A$ and disease $B$. The system of differential equations for this flow-diagram are represented in System LABEL:eqn:SIRS with functions for $\beta_{i}$ in Table 1 and $m_{i}$ in Equation (2). Solid arrows represent flow of individuals. Dashed lines represent flow of information.
We incorporate the effect of risk response on transmission such that as prevalence increases (decreases) transmission declines (rises), representing human response to risk. We model this behavioral response, $m_{i}$, as
$m_{i}=\exp\left(-k\widetilde{I}_{i}\right),$ (2)
which represents the effect of response (e.g., decline in contact rate) on the transmission rate of $\beta_{i}$. Here, $m_{i}$ is formulated such that increases in prevalence ($I$), tracked through a lagged compartment of the infectious population, $\widetilde{I}_{i}$, lead to a decline in $m_{i}$. As prevalence $I_{i}$ increases, perceived risk of $\widetilde{I}_{i}$ (modeled by an exponential delay of prevalence) also increases. Higher perceived risk leads to stronger reaction, decreasing $m_{i}$, and consequently decreasing $\beta_{i}$.
2.1 Transmission rate
We consider three scenarios that differ with respect to the level of behavioral spillover. Mathematically, this leads to differences in the multipliers $m_{A}$ and $m_{B}$ in the formulation of the transmission rate, denoted by $\beta_{i}(\cdot)$, in the SIRS model given in System (LABEL:eqn:SIRS). The specific system of equations for each scenario, with the relevant formulation of transmission, is included in Appendix A1.1 for the reader’s reference.
2.1.1 Two independent diseases
Scenario 1: No spillover.
For the case of two pathogens that are not behaviorally coupled and with no cross-immunity, we assume that the transmission rate of each pathogen is formulated independently as
$\beta_{i}=m_{i}\beta_{0,i},$
where $\beta_{0,i}$ is the infectivity of disease $i$ and $m_{i}$ the multiplier of disease $i$. In this case, disease $A$ and disease $B$ are spreading independently.
2.1.2 Two behaviorally-coupled diseases
Here, we retain the assumption of no cross-immunity between the diseases, but include behavioral interdependencies, where NPIs implemented to contain one disease can influence the spread of the other one. Specifically, we connect the two diseases $A$ and $B$ via their behavioral mechanism. We consider two scenarios: first, perfect spillover, where the impact of behavioral responses to one disease perfectly affects the spread of the other one, possibly due to very similar transmission mechanisms; and second, a more general context where the impact of behavioral response targeting one disease is only partially effective on the other one.
Scenario 2: Perfect spillover.
As a first scenario, we assume the simple case of perfect spillover. This represents a situation where a proportional decline in transmission of disease $A$, due to behavioral response, results in the same proportional decline in the transmission of disease $B$, and vice versa. This happens when the two diseases transmit in an identical manner. Behavioral response to one can alter the spread of the other, such that transmission of each disease is affected by both $m_{A}$ and $m_{B}$. Thus, we modify the transmission rate of disease $i$, $\beta_{i}$, to be
$\beta_{i}$ $=\beta_{0,i}m_{A}m_{B},$
$=\beta_{0,i}\exp\left(-k\left(\widetilde{I}_{A}+\widetilde{I}_{B}\right)\right).$
Scenario 3: Imperfect spillover.
For a more general description, we consider imperfect spillover, where a decline in transmission of one disease results in an $s$ fractional decline in transmission of the other disease. For example, we express $\beta_{A}$ as
$\beta_{A}$ $=m_{A}\left(1-s\left(1-m_{B}\right)\right)\beta_{0,A}.$
Formally, the effect of risk response to the prevalence of $B$ on the spread of disease $A$ is $\left(1-s\left(1-m_{B}\right)\right)$. We formulate $\beta_{B}$ in a similar manner (see Table 1).
It is important to note that this formulation encapsulates both no spillover (independent diseases) and perfect spillover. For perfect spillover ($s=1$), the overall effect of the response to $B$ on the spread of disease $A$ is $m_{B}$, and for no spillover ($s=0$), the overall effect of response to disease $B$ on the spread of $A$ is constant at one. Table 1 summarizes all the formulations for $\beta_{i}(\cdot)$ with $i\in\{A,B\}$. We note that in model formulation and analytical results, we present no spillover ($s=0$) and perfect spillover ($s=1$) before discussing imperfect spillover ($0<s<1$), while for numerical results, we present results ordered according to increasing spillover.
Table 1: Formulations of transmission rate, $\beta_{i}(\cdot)$ for two diseases, AA and $B$, under no spillover, perfect spillover and imperfect spillover.
Transmission rate of disease AA Transmission rate of disease $B$
Independent diseases No spillover ($s=0$) $\begin{aligned} \beta_{A}&=m_{A}\beta_{0,A}\\ &=e^{-k\widetilde{I}_{A}}\beta_{0,A}\end{aligned}$ $\begin{aligned} \beta_{B}&=m_{B}\beta_{0,B}\\ &=e^{-k\widetilde{I}_{B}}\beta_{0,B}\end{aligned}$
Behaviorally-coupled diseases Perfect spillover ($s=1$) $\begin{aligned} \beta_{A}&=m_{A}m_{B}\beta_{0,A}\\ &=e^{-k\widetilde{I}_{A}}e^{-k\widetilde{I}_{B}}\beta_{0,A}\end{aligned}$ $\begin{aligned} \beta_{B}&=m_{A}m_{B}\beta_{0,B}\\ &=e^{-k\widetilde{I}_{A}}e^{-k\widetilde{I}_{B}}\beta_{0,B}\end{aligned}$
Behaviorally-coupled diseases Imperfect spillover ($0<s<1$) $\begin{aligned} \beta_{A}&=m_{A}\big(1-s(1-m_{B})\big)\beta_{0,A}\\ &=e^{-k\widetilde{I}_{A}}\big(1-s(1-e^{-k\widetilde{I}_{B}})\big)\beta_{0,B}\end{aligned}$ $\begin{aligned} \beta_{B}&=\big(1-s(1-m_{A})\big)m_{B}\beta_{0,A}\\ &=\big(1-s(1-e^{-k\widetilde{I}_{A}})\big)e^{-k\widetilde{I}_{B}}\beta_{0,B}\end{aligned}$
3 Results
We begin by numerically simulating the dynamics of our system with a range of spillover – none, perfect and imperfect (Section 3.1). We choose parameters from the literature on COVID-19 and influenza to highlight our application of this model to concrete scenarios. The parameter values used for analysis and simulation of the two diseases ($A$ and $B$) are given in Table 2. Following our numerical examination, we determine the existence and stability of equilibria of the model under each spillover scenario (Section 3.2) and assess the structural and practical identifiability of the model (Section 3.3).
3.1 Numerical results
To examine the dynamics of our behaviorally-coupled system, we numerically simulate our system in MATLAB R2021a [28] using ode45 and observe the dynamics over short (one year) and long (multiple years) time scales. Figure 3 shows results from nine simulation experiments which vary the level of spillover and the basic reproduction number of disease $B$ (with the basic reproduction number of disease $A$ fixed at 3). A helpful quantity for explaining the behavior in the system is the effective reproduction number, $\mathcal{R}_{e}$, which tells if the disease is expanding or contracting at the current time. For our two disease system, we define the effective reproduction number of disease $i$, $\mathcal{R}_{e,i}$, as
$\mathcal{R}_{e,i}$ $=\beta_{i}S_{i}\tau_{I},$
$=m_{i}(1-s(1-m_{j}))\beta_{0,i}S_{i}\tau_{I},$
$=m_{i}(1-s(1-m_{j}))S_{i}R_{0,i},$
where $m_{j}$ denotes dependence on disease $j$ ($j\neq i$). Without loss of generality, we consider the scenario where $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$. Here, the spillover level, $s$, falls between zero and one, where zero indicates no spillover (independent diseases) and one indicates perfect spillover.
Table 2: Symbol, description and values for parameter and initial conditions. The values are chosen to represent spillover between two similar diseases (e.g., COVID-19 and Influenza).
Symbol Description Value Source
$\mathcal{R}_{0,A}$ basic reproduction number of disease AA 3 [13]
$\mathcal{R}_{0,B}$ basic reproduction number of disease $B$ [1,3] [49]
$\beta_{0,A}$ infectivity rate of disease AA $\frac{\mathcal{R}_{0,A}}{\tau_{I}}$ [13, 8]
$\beta_{0,B}$ infectivity rate of disease $B$ $\frac{\mathcal{R}_{0,B}}{\tau_{I}}$ [49, 8]
$\tau_{I}$ infection period 7 days [8]
$\tau_{P}$ delay to adjust risk perception 30 days [57]
$\tau_{R}$ immunity period 100 days [63]
$k$ sensitivity to risk 100 [23]
$s$ spillover constant 0 varied in [0,1]
$S_{i}(0)$ initial proportion of disease-$i$ susceptible individuals 0.9999 assumed
$I_{i}(0)$ initial proportion of disease-$i$ infectious individuals 0.0001 assumed
$R_{i}(0)$ initial proportion of disease-$i$ recovered individuals 0 assumed
$\widetilde{I}_{i}(0)$ initial proportion of disease-$i$ perceived infectious individuals 0 assumed
Without spillover (under our assumption $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$), there are two regimes of behavior: (i) diseases $A$ and $B$ co-exist but disease $A$ prevalence always remains above disease $B$ prevalence and (ii) diseases $A$ and $B$ co-exist but the diseases alternate whose prevalence is higher. As depicted in the left column of Figure 3 (Panels 3(a), 3(d), 3(g)), both diseases persist as the diseases are spreading independently and (by assumption) the basic reproduction number of each disease is above one. As the diseases are not interacting in any way, the endemic level of the disease is dictated solely by its basic reproduction number. Furthermore, the dynamics are qualitatively similar for the two diseases and are quantitatively identical when the diseases have the same traits. As cases grow for disease $i$, individuals respond, leading to reductions in the transmission rate and a drop in $\mathcal{R}_{e,i}$. Ultimately, $\mathcal{R}_{e,i}$ settles to one such that cases reach a non-zero equilibrium (under our assumption $\mathcal{R}_{0,i}>1$). As the basic reproduction number of disease $B$ is lowered, the initial peak occurs later and reaches a lower level (Figure 3adg, red lines as $\mathcal{R}_{0,B}$ is lowered from 2.9 to 2 to 1.3). This is because the natural disease infectivity is lower, reducing the transmission rate even without behavioral response. Thus, for a range of differences in $\mathcal{R}_{0,i}$ for the two diseases – even in the absence of interdependent behavioral feedback (i.e., diseases spread entirely independently) – the diseases may alternate in prevalence levels, as seen in Figure 3ad, an example of regime (ii). When $\mathcal{R}_{0,A}>>\mathcal{R}_{0,B}$ (Figure 3g), while the disease co-exist, disease $A$ always has higher prevalence, an example of regime (i).
Figure 3: Dynamics with three levels of spillover and three values of the basic reproduction number of disease $B$. For (a), (d) and (g), there is no spillover, $s=0$, which corresponds to the two diseases spreading independently (i.e., the curve is identical to that of the single disease spreading in the population); for (b), (e), (h), there is imperfect spillover, $s=0.5$; and for (c), (f), (i), there is perfect spillover, $s=1$. For (a)-(c), $\mathcal{R}_{0,B}=2.9$; for (d)-(f), $\mathcal{R}_{0,B}=2$; and for (g)-(i) $\mathcal{R}_{0,B}=1.3$). Other parameters as in Table 2 with $\mathcal{R}_{0,A}=3$.
In the case of behavioral spillover, the dynamics become more interesting. The middle and right columns of Figure 3 show results from the scenarios of imperfect and perfect spillover, respectively. With our assumption of a higher basic reproduction number for disease AA ($\mathcal{R}_{0,A}>\mathcal{R}_{0,B}$), cases initially grow for disease $A$ faster than for disease $B$ and thus, people react to disease $A$ sooner than to disease $B$. As a result, the $\mathcal{R}_{e,A}$ falls, but depending on $\mathcal{R}_{0,B}$ and the level of spillover, this may provide the opportunity for disease $B$ to flourish. Here, there are now three regimes of behavior: (i) diseases $A$ and $B$ co-exist and disease $A$ prevalence always remains above disease $B$ prevalence, (ii) diseases $A$ and $B$ co-exist but alternate whose prevalence is higher, and (iii) only disease $A$ persists, as disease $B$ is unable to persist despite a basic reproduction number above one.
When there is behavioral spillover, disease $A$ and disease $B$ impact each other’s spread. This alters the region of co-existence of both diseases. Under our assumption of $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$, disease $A$ always persists but the region where disease $B$ also persists is determined by the combination of spillover ($0<s\leq 1$) and the transmission capacity of disease $B$, as determined by $\mathcal{R}_{0,B}$. The impact of disease $A$ on disease $B$ potentially leads to its extinction even when $\mathcal{R}_{0,B}>1$. For example, regime (iii) is seen in Figure 3fhi.
We conduct a more systematic analysis, varying values of $\mathcal{R}_{0,B}$ and $s$ in small increments, and examine the final simulation results for outcome regimes. Specifically, $\mathcal{R}_{0,B}$ is iterated between 1 and 3 (with $\mathcal{R}_{0,A}=3$) and $s$ is iterated between 0 (no spillover) and 1 (perfect spillover). Figure 4 shows simulated outcomes over the first year of an epidemic for various combinations of the two parameter values. As seen in Figure 4a, as $\mathcal{R}_{0,B}$ increases, a larger amount of spillover (larger $s$) is required for exclusion of disease $B$. In Figure 4b, we quantify the percentage of time in the first year where the prevalence of disease $B$ exceeds that of disease $A$. However, when disease $B$ prevalence exceeds disease $A$, it may not be by much. Thus, in Figures 4c and d we quantify the cumulative amount across one year by which one disease exceeds the other. The specified nine points marked in Figure 4 correspond to the scenarios presented in Figure 3. Considering the entire range of possible spillover and basic reproduction numbers of disease $B$ above one but below that of disease AA ($\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$), co-existence is always possible for low spillover, but requires a large basic reproduction number for strain $B$ with high spillover (Figure 4a). Furthermore, only for high enough $\mathcal{R}_{0,B}$ and low enough spillover can strain $B$ rise in prevalence above strain $A$ at any point (Figure 4b) and the level by which it exceeds disease $A$ may be quite low (Figure 4d). Our results on the impact of behavioral spillover are robust to another choice of the behavioral feedback, as demonstrated by the use of a fractional rather than exponential formulation (see [38] for discussion of various functional formulations for behavioral feedback). In particular, we reproduce Figs. 3 and 4 in Section A2 with a discussion of the fractional formulation.
(a) Co-existence vs. exclusion (simulation)
(b) Time disease $B$ exceeds disease AA
(c) Amount disease $A$ exceeds disease $B$
(d) Amount disease $B$ exceeds disease AA
Figure 4: Persistence and superiority of diseases across one year as computed through numerical simulation. (a) Persistence of both diseases (purple) or only disease AA (red), which has the higher basic reproduction number. (b) Percentage of the first year that disease $B$ prevalence is above disease AA prevalence. For $\mathcal{R}_{0,A}=\mathcal{R}_{0,B}=3$ the diseases produce independent, identical outbreaks so neither disease exceeds the other one (white line at the tope). White space at the bottom right corresponds to when there is persistence of only disease $A$. (c) Cumulative amount during the first year that disease $A$ exceeds disease $B$. (d) Cumulative amount during the first year that disease $B$ exceeds disease $A$. Note the different color scaling in panels (b)-(d). Dots correspond to combination of spillover ($s$) and basic reproduction number of disease $B$ ($\mathcal{R}_{0,B}$) used for plots in Figure 3.
3.1.1 Approximation of spillover
To better understand the role of spillover in co-existence of both diseases, we assess our analytical conditions of stability equilibria (see Section 3.2 for their calculation) as well as provide an approximation of the minimum spillover required in order for disease $A$ to exclude disease $B$. See Theorem 1 below. The combination of the level of spillover and the basic reproduction number for disease $B$ (under our assumption that $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$) in order for co-existence of both diseases is seen in Figure 5. Compared to our numerical simulations in Figure 4a and our numerically-evaluated analytical conditions in Figure 5b (which are visually nearly identical), our approximation (shown in Figure 5a) under-estimates the parameter regime for which there is co-existence, i.e., co-existence is still possible for larger spillover and smaller basic reproduction number than the approximation suggests. Furthermore, there are slight differences between our simulation based results (Figure 4a) and our numerically-calculated analytical results (Figure 5b), such as for low $\mathcal{R}_{0,B}$ and low $s$, as we assess persistence of disease $B$ in the simulation results as falling below the threshold of $10^{-4}$ at one year.
(a) Co-existence vs. exclusion (approximation)
(b) Co-existence vs. exclusion (analytical)
Figure 5: Analytical approximation and analytical solution of regions of disease persistence. (a) Approximated threshold for co-existence of both diseases as determined in Theorem 1. (b) Analytical solution (numerically computed) of persistence of both diseases (purple) or only disease AA (red). Recall that disease AA has, by assumption, the higher basic reproduction number ($\mathcal{R}_{0,A}=3$). Dots correspond to combination of spillover ($s$) and basic reproduction number of disease $B$ ($\mathcal{R}_{0,B}$) used for plots in Figure 3.
Theorem 1.
Approximated threshold of spillover for co-existence of diseases. Without loss of generality, assuming $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$, the spillover fraction, $s$, needs to be above $s_{\text{threshold}}$, i.e.,
$s_{\text{threshold}}=\frac{1-\frac{1}{\mathcal{R}_{0,B}}}{1-\frac{1}{\mathcal{R}_{0,A}}}<s<1,$
for the possibility of exclusion of strain $B$.
To provide an intuition about the threshold, consider ${\mathcal{R}_{0,A}}=3$ and ${\mathcal{R}_{0,B}}=1.3$ (as in the bottom row of Figure 3), which may resemble the case of SARS-CoV-2 and an influenza-like virus. In this case, $s_{threshold}\approx 0.35$, such that for higher values of spillover ($s>0.35$), there is the possibility of complete suppression of disease $B$, the influenza-like virus. We observe this suppression when $s=0.5$ in Figure 3, bottom row, middle column, where ${\mathcal{R}_{0,A}}=3$ and ${\mathcal{R}_{0,B}}=1.3$.
3.2 Equilibria and their stability
Even though epidemics exhibit complex dynamics and such systems may rarely remain at equilibrium due to external perturbations or seasonal forcing, equilibrium analysis provides insight into the underlying mechanisms of disease spread and helps guide intuition about the system’s dynamics.
As suggested by Figure 3, the system exhibits multiple equilibria. In this section, we establish conditions for the existence, uniqueness, and stability of all possible equilibria for the system under the cases when $s=0$ (no spillover), $s=1$ (perfect spillover), and $s\in(0,1)$ (imperfect spillover). Theorem 2 establishes the existence of the equilibria for the three scenarios. Theorem 3 describes the stability of the disease-free equilibrium (DFE) and its association to the basic reproduction numbers of each disease. Theorems 4 and 5 establish the stability of the boundary equilibria and endemic equilibrium, respectively. The proofs for each theorem are found in the appendix. We summarize the existence and uniqueness conditions as well as the stability conditions for no spillover in Table 3, for perfect spillover in Table 4, and for imperfect spillover in Table 5.
Theorem 2.
Existence of equilibria for the three scenarios
When considering two diseases ($A$ and $B$), given the biologically relevant (positive) parameters and realistic (non-negative and bounded) initial conditions, each of the three scenarios (no spillover, perfect spillover, and imperfect spillover) has four equilibria, given below, written with the compartments in the following order
$\left(S_{A},I_{A},R_{A},\widetilde{I}_{A},S_{B},I_{B},R_{B},\widetilde{I}_{B}\right).$
Solutions for each of the three scenarios – no spillover, perfect spillover, and imperfect spillover – are denoted by $\mathcal{E}^{*}$, $\mathcal$$\hat${E}}, and $\mathcal$$\bar${E}}, respectively. We denote the equilibrium values for the infectious populations as $I_{i}=I_{i}^{*}$ for $\mathcal{E}^{*}$; $I_{i}=\hat{I}_{i}$ for $\mathcal$$\hat${E}}; and $I_{i}=\bar{I}_{i}$ for $\mathcal$$\bar${E}}. Here, $i\in\{A,~B,~A_{E},~B_{E}\}$, with $I_{A}$, $I_{B}$ representing the equilibrium values where only disease $A$ or $B$ are endemic, respectively, and $I_{A_{E}}$, $I_{B_{E}}$ representing the equilibrium values where both diseases are endemic.
  1. 1.
    The disease-free equilibrium for each scenario ($\mathcal{E}_{0}^{*}$, $\hat{\mathcal{E}}_{0}$, $\bar{\mathcal{E}}_{0}$, respectively)
    $(1,0,0,0,1,0,0,0)$
    always exists.
  2. 2.
    The disease-B-free and disease-A-endemic equilibrium (disease-$A$ “boundary equilibrium”) for each scenario ($\mathcal{E}_{A}^{*}$, $\hat{\mathcal{E}}_{A}$, $\bar{\mathcal{E}}_{A}$, respectively)
    $\left(1-I_{A}-\dfrac{\tau_{R}}{\tau_{I}}I_{A},I_{A},\dfrac{\tau_{R}}{\tau_{I}}I_{A},I_{A},1,0,0,0\right)$
    exists if and only if $\beta_{0,A}\tau_{I}>1$. Further, it follows that $\beta_{0,A}\tau_{I}>\exp(kI^{*}_{A})$, $\beta_{0,A}\tau_{I}>\exp(k\hat{I}_{A})$, and $\beta_{0,A}\tau_{I}>\exp(k\bar{I}_{A})$ when these equilibria exist.
  3. 3.
    The disease-A-free and disease-B-endemic equilibrium (disease-$B$ “boundary equilibrium”) for each scenario ($\mathcal{E}_{B}^{*}$, $\hat{\mathcal{E}}_{B}$, $\bar{\mathcal{E}}_{B}$, respectively)
    $\left(1,0,0,0,1-I_{B}-\dfrac{\tau_{R}}{\tau_{I}}I_{B},I_{B},\dfrac{\tau_{R}}{\tau_{I}}I_{B},I_{B}\right)$
    exists if and only if $\beta_{0,B}\tau_{I}>1$. Further, it follows that $\beta_{0,B}\tau_{I}>\exp(kI^{*}_{B})$, $\beta_{0,B}\tau_{I}>\exp(k\hat{I}_{B})$, and $\beta_{0,B}\tau_{I}>exp(k\bar{I}_{B})$ when these equilibria exist.
  4. 4.
    The endemic equilibrium for each scenario ($\mathcal{E}_{A,B}^{*}$, $\hat{\mathcal{E}}_{A,B}$, $\bar{\mathcal{E}}_{A,B}$, respectively)
    $\left(1-{I}_{A_{E}}-\dfrac{\tau_{R}}{\tau_{I}}{I}_{A_{E}},{I}_{A_{E}},\dfrac{\tau_{R}}{\tau_{I}}{I}_{A_{E}},{I}_{A_{E}},1-{I}_{B_{E}}-\dfrac{\tau_{R}}{\tau_{I}}{I}_{B_{E}},{I}_{B_{E}},\dfrac{\tau_{R}}{\tau_{I}}{I}_{B_{E}},{I}_{B_{E}}\right),$
    exists if and only if
    1. (a)
      $\beta_{0,A}\tau_{I}>1$ and $\beta_{0,B}\tau_{I}>1$ for $\mathcal{E}_{A,B}^{*}$;
    2. (b)
      $\beta_{0,A}\tau_{I}>e^{k\hat{I}_{B_{E}}}$ and $\beta_{0,B}\tau_{I}>e^{k\hat{I}_{A_{E}}}$ for $\hat{\mathcal{E}}_{A,B}$; and
    3. (c)
      $\beta_{0,A}\tau_{I}>\dfrac{\exp(k\bar{I}_{B_{E}})}{\exp(k\bar{I}_{B_{E}})-s(\exp(k\bar{I}_{B_{E}})-1)}$ and $\beta_{0,B}\tau_{I}>\dfrac{\exp(k\bar{I}_{A_{E}})}{\exp(k\bar{I}_{A_{E}})-s(\exp(k\bar{I}_{A_{E}})-1)}$ for $\bar{\mathcal{E}}_{A,B}$.
From Theorem 2, we observe that all disease-specific endemic boundary equilibrium values are equal (along with the endemic equilibrium values for the case of no spillover ($s=0$)), regardless of spillover; that is, $I^{*}_{A}=\hat{I}_{A}=\bar{I}_{A}=I^{*}_{A_{E}}$, and $I^{*}_{B}=\hat{I}_{B}=\bar{I}_{B}=I^{*}_{B_{E}}$.
Table 3: Conditions for existences and uniqueness as well as stability of equilibria with no spillover ($s=0$).
No spillover ($s=0$)
$\mathcal{E}_{0}^{*}$ $\mathcal{E}_{A}^{*}$ $\mathcal{E}_{B}^{*}$ $\mathcal{E}_{A,B}^{*}$
$I_{A}=0$, $I_{A}=I^{*}_{A}>0$, $I_{A}=0$, $I_{A}=I^{*}_{A_{E}}>0$,
$I_{B}=0$ $I_{B}=0$ $I_{B}=I^{*}_{B}>0$ $I_{B}=I^{*}_{B_{E}}>0$
Existence & uniqueness Always exists $\mathcal{R}_{0,A}>1$ $\mathcal{R}_{0,B}>1$ $\mathcal{R}_{0,A}>1$, $\mathcal{R}_{0,B}>1$,
Stability $\mathcal{R}_{0,A}<1$, $\mathcal{R}_{0,B}<1$ $\mathcal{R}_{0,A}>1$, $\mathcal{R}_{0,B}<1$ $\mathcal{R}_{0,A}<1$, $\mathcal{R}_{0,B}>1$ $\mathcal{R}_{0,A}>1$ $\mathcal{R}_{0,B}>1$
For Theorems 3 - 5, regarding stability of equilibria, we simplify our system by noting that $R_{i}=1-S_{i}-I_{i}$, as $R_{i}$ can be determined by $S_{i}$ and $I_{i}$, and omit $R_{i}$ when carrying out the analysis. To this end, we make the substitution for $R_{i}$ in each of the three scenarios and consider only $\dfrac{dS_{i}}{dt}$, $\dfrac{dI_{i}}{dt}$, and $\dfrac{d\widetilde{I}_{i}}{dt}$ in each, reducing to the six-dimensional subsystem.
Theorem 3.
Stability of disease-free equilibrium (DFE) $\mathcal{E}_{0}^{*}$, $\hat{\mathcal{E}}_{0}$, $\bar{\mathcal{E}}_{0}$
Given biologically relevant parameters and realistic initial conditions, the disease-free equilibrium is locally asymptotically stable for all three scenarios if $\mathcal{R}_{0,A}:=\beta_{0,A}\tau_{I}<1$ and $\mathcal{R}_{0,B}:=\beta_{0,B}\tau_{I}<1$.
Table 4: Conditions for existences and uniqueness as well as stability for equilibria with perfect spillover ($s=1$). Note that the condition for the stability of $\bar{\mathcal{E}}_{A,B}$ is a conjecture.
Perfect Spillover ($s=1$)
$\hat{\mathcal{E}}_{0}$ $\hat{\mathcal{E}}_{A}$ $\hat{\mathcal{E}}_{B}$ $\hat{\mathcal{E}}_{A,B}$
$I_{A}=0$, $I_{A}=\hat{I}_{A}>0$, $I_{A}=0$, $I_{A}=\hat{I}_{A_{E}}>0$,
$I_{B}=0$ $I_{B}=0$ $I_{B}=\hat{I}_{B}>0$ $I_{B}=\hat{I}_{B_{E}}>0$
Existence & uniqueness Always exists $\mathcal{R}_{0,A}>1$ $\mathcal{R}_{0,B}>1$ $\mathcal{R}_{0,A}>e^{k\hat{I}_{B}}$, $\mathcal{R}_{0,B}>e^{k\hat{I}_{A}}$
Stability $\mathcal{R}_{0,A}<1$, $\mathcal{R}_{0,B}<1$ $\mathcal{R}_{0,A}>1$, $\mathcal{R}_{0,B}<e^{k\hat{I}_{A}}$, ($\implies\mathcal{R}_{0,A}>\mathcal{R}_{0,B}$) $\mathcal{R}_{0,A}<e^{k\hat{I}_{B}}$, $\mathcal{R}_{0,B}>1$, ($\implies\mathcal{R}_{0,B}>\mathcal{R}_{0,A}$) $\mathcal{R}_{0,A}>e^{k\hat{I}_{B}}$, $\mathcal{R}_{0,B}>e^{k\hat{I}_{A}}$
Theorem 4.
Stability of boundary equilibria $\mathcal{E}_{A}^{*},\hat{\mathcal{E}}_{A},\bar{\mathcal{E}}_{A}$ and $\mathcal{E}_{B}^{*},\hat{\mathcal{E}}_{B},\bar{\mathcal{E}}_{B}$
Given biologically relevant parameters and realistic initial conditions, for each scenario, we have the following conditions for stability of the boundary equilibria:
  1. 1.
    $\mathcal{E}_{A}^{*}$ will be locally asymptotically stable if $\mathcal{R}_{0,A}>1$ and $\mathcal{R}_{0,B}<1$;
  2. 2.
    $\mathcal{E}_{B}^{*}$ will be locally asymptotically stable if $\mathcal{R}_{0,B}>1$ and $\mathcal{R}_{0,A}<1$;
  3. 3.
    $\hat{\mathcal{E}}_{A}$ will be locally asymptotically stable if $\mathcal{R}_{0,A}>1$ and $\mathcal{R}_{0,B}<\exp(k\hat{I}_{A})$, which, along with the existence condition, implies $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}$;
  4. 4.
    $\hat{\mathcal{E}}_{B}$ will be locally asymptotically stable if $\mathcal{R}_{0,B}>1$ and $\mathcal{R}_{0,A}<\exp(k\hat{I}_{B})$, which, along with the existence condition, implies $\mathcal{R}_{0,B}>\mathcal{R}_{0,A}$;
  5. 5.
    $\bar{\mathcal{E}}_{A}$ will be locally asymptotically stable if $\mathcal{R}_{0,A}>1$ and $\mathcal{R}_{0,B}<\dfrac{\exp(k\bar{I}_{A})}{\exp(k\bar{I}_{A})-s(\exp(k\bar{I}_{A})-1)}$, which, along with the existence condition, implies $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}$; and
  6. 6.
    $\bar{\mathcal{E}}_{B}$ will be locally asymptotically stable if $\mathcal{R}_{0,B}>1$ and $\mathcal{R}_{0,A}<\dfrac{\exp(k\bar{I}_{B})}{\exp(k\bar{I}_{B})-s(\exp(k\bar{I}_{B})-1)}$, which, along with the existence condition, implies $\mathcal{R}_{0,B}>\mathcal{R}_{0,A}$.
Note that, in the case of spillover (either perfect or imperfect), we can have both $\mathcal{R}_{0,A},\mathcal{R}_{0,B}>1$, a condition which typically results in long-term establishment of both diseases, with the boundary equilibria still stable. In other words, the disease with the lower basic reproduction number will not be successful in persisting and will eventually become extinct, as long as its $\mathcal{R}_{0}$ is smaller than the respective endemic equilibrium existence condition for $0<s\leq 1$, i.e.,
$\frac{e^{k\bar{I}_{B}}}{e^{k\bar{I}_{B}}-s(e^{k\bar{I}_{B}}-1)}>\mathcal{R}_{0,A}>1\qquad\text{or}\qquad\frac{e^{k\bar{I}_{A}}}{e^{k\bar{I}_{A}}-s(e^{k\bar{I}_{A}})-1)}>\mathcal{R}_{0,B}>1.$
These conditions being strictly greater than one follow from the denominator always being smaller than the numerator for any amount of spillover, $0<s\leq 1$. For example, the boundary equilibrium with only disease $A$ for imperfect spillover, $\bar{\mathcal{E}}_{A}$, will be locally asymptotically stable, such that disease $B$ will return to its disease-free state after an initial outbreak, if $\mathcal{R}_{0,B}<\dfrac{e^{k\bar{I}_{A}}}{e^{k\bar{I}_{A}}-s(e^{k\bar{I}_{A}})-1)}$, even though both $\mathcal{R}_{0,A}>1$ and $\mathcal{R}_{0,B}>1$. Note that when this condition is met, it also holds that $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}$. We see this behavior exhibited in Figure 3h for $s=0.5$ with $\mathcal{R}_{0,B}=1.3$. Similar behavior occurs with perfect spillover as in Figure 3f for $s=1$ with $\mathcal{R}_{0,B}=2$; and in Figure 3i for $s=1$ with $\mathcal{R}_{0,B}=1.3$.
Theorem 5.
Stability of $\mathcal{E}^{*}_{A,B}$
Given biologically relevant parameters and realistic initial conditions, $\mathcal{E}^{*}_{A,B}$ will be locally asymptotically stable if $\mathcal{R}_{0,A}>1$ and $\mathcal{R}_{0,B}>1$.
Theorem 5 only addresses the stability of the endemic equilibrium, $\mathcal{E}^{*}_{A,B}$, for no spillover. In addition, we present conjectures for the stability of the endemic equilibria in the case of perfect spillover (Table 4) and imperfect spillover (Table 5).
Table 5: Conditions for existences and uniqueness as well as stability for equilibria with imperfect spillover ($0<s<1$). Note that the condition for the stability of $\bar{\mathcal{E}}_{A,B}$ is a conjecture.
Imperfect Spillover ($0<s<1$)
$\bar{\mathcal{E}}_{0}$ $\bar{\mathcal{E}}_{A}$ $\bar{\mathcal{E}}_{B}$ $\bar{\mathcal{E}}_{A,B}$
$I_{A}=0$, $I_{A}=\bar{I}_{A}>0$, $I_{A}=0$, $I_{A}=\bar{I}_{A_{E}}>0$,
$I_{B}=0$ $I_{B}=0$ $I_{B}=\bar{I}_{B}>0$ $I_{B}=\bar{I}_{B_{E}}>0$
Existence & uniqueness Always exists $\mathcal{R}_{0,A}>1$ $\mathcal{R}_{0,B}>1$ $\mathcal{R}_{0,A}>\frac{e^{k\bar{I}_{B}}}{e^{k\bar{I}_{B}}-s(e^{k\bar{I}_{B}}-1)}$, $\mathcal{R}_{0,B}>\frac{e^{k\bar{I}_{A}}}{e^{k\bar{I}_{A}}-s(e^{k\bar{I}_{A}})-1)}$
Stability $\mathcal{R}_{0,A}<1$, $\mathcal{R}_{0,B}<1$ $\mathcal{R}_{0,A}>1$, $\mathcal{R}_{0,B}<\frac{e^{k\bar{I}_{A}}}{e^{k\bar{I}_{A}}-s(e^{k\bar{I}_{A}}-1)}$, ($\implies\mathcal{R}_{0,A}>\mathcal{R}_{0,B}$) $\mathcal{R}_{0,A}<\frac{e^{k\bar{I}_{B}}}{e^{k\bar{I}_{B}}-s(e^{k\bar{I}_{B}}-1)}$, $\mathcal{R}_{0,B}>1$, ($\implies\mathcal{R}_{0,B}>\mathcal{R}_{0,A}$) $\mathcal{R}_{0,A}>\frac{e^{k\bar{I}_{B}}}{e^{k\bar{I}_{B}}-s(e^{k\bar{I}_{B}}-1)}$, $\mathcal{R}_{0,B}>\frac{e^{k\bar{I}_{A}}}{e^{k\bar{I}_{A}}-s(e^{k\bar{I}_{A}}-1)}$
3.3 Identifiability
In order to determine whether the parameter values can be uniquely estimated from the model formulation, we conduct an identifiability analysis. First, we perform structural identifiability analysis using the software STRIKE-GOLDD 4.2 to determine whether the parameters are globally identifiable, locally identifiable, or unidentifiable [15] (Section 3.3.1). For scenarios that are structurally identical, we consider a practical identifiability analysis using the Monte Carlo approach to establish which parameters are strongly identifiable, weakly identifiable, or unidentifiable (Section 3.3.2).
3.3.1 Structural Identifiability
Mathematically, we say that a model is globally structurally identifiable for a set of parameters $\mathbf{p_{1}}$ for output $\mathbf{y}(t)$, if for every parameter set $\mathbf{p_{2}}$, the relationship $y(t,\mathbf{p_{1}})=y(t,\mathbf{p_{2}})\hskip 5.69046pt\text{implies}\hskip 5.69046pt\mathbf{p_{1}}=\mathbf{p_{2}}.$ If there is a neighborhood in which the equation above holds true, we say the model is locally identifiable. We consider three outputs: prevalence (level of infectious classes), incidence (entrance into the infectious classes), and ‘recognized prevalence’ (a scaled version of perceived prevalence, which relates to the infectious level following a delay). Here we assess parameters: $\beta_{A},\beta_{B},k,\tau_{R},\tau_{I},\tau_{P}$ for prevalence ($y_{1}(t)=I_{A}(t)$ and $y_{2}(t)=I_{B}(t)$), incidence ($y_{1}(t)=\beta_{A}S_{A}(t)I_{A}(t)$ and $y_{2}(t)=\beta_{B}S_{B}(t)I_{B}(t)$), and recognized prevalence ($y_{1}(t)=K_{A}\widetilde{I}_{A}(t)$ and $y_{2}(t)=K_{B}\widetilde{I}_{B}(t)$, where $K_{A}$ and $K_{B}$ are a fraction of perceived prevalence).
We employ STRIKE-GOLDD 4.2 which uses Lie derivatives to obtain information about the state variables of nonlinear systems from the output measures [69]. Since our models contain an exponential term, we cannot use the traditional software such as DAISY or Julia as these tools are dependent on the differential algebra methodology [17, 5]. STRIKE-GOLDD 4.2 is implemented in MATLAB R2021a.
Table 6: The identifiability results using the three formulations described in Table 1 using the toolbox STRIKE-GOLDD 4.2. Note that * means MATLAB ran out of memory to complete the simulation.
Case Outputs Structural Identifiability Results
Independent diseases No spillover ($s=0$) $\begin{aligned} y_{1}(t)&=I_{A}(t)\\ y_{2}(t)&=I_{B}(t)\end{aligned}$ $\beta_{A},\beta_{B},k,\tau_{R},\tau_{I},\tau_{P}$ are all locally identifiable.
$\begin{aligned} y_{1}(t)&=\beta_{A}S_{A}(t)I_{A}(t)\\ y_{2}(t)&=\beta_{B}S_{B}(t)I_{B}(t)\end{aligned}$ $\beta_{A},\beta_{B},k,\tau_{R},\tau_{I},\tau_{P}$ are all locally identifiable.
$\begin{aligned} y_{1}(t)&=K_{A}\widetilde{I}_{A}(t)\\ y_{2}(t)&=K_{B}\widetilde{I}_{B}(t)\end{aligned}$ $\beta_{A},\beta_{B},k,K_{A},K_{B},\tau_{R},\tau_{I},\tau_{P}$ are all locally identifiable.
Behaviorally-coupled diseases Perfect spillover ($s=1$) $\begin{aligned} y_{1}(t)&=I_{A}(t)\\ y_{2}(t)&=I_{B}(t)\end{aligned}$ $\beta_{A},\beta_{B},k,\tau_{R},\tau_{I},\tau_{P}$ are all locally identifiable. $\widetilde{I}_{A}(0)$ and $\widetilde{I}_{B}(0)$ are unidentifiable.
$\begin{aligned} y_{1}(t)&=\beta_{A}S_{A}(t)I_{A}(t)\\ y_{2}(t&)=\beta_{B}S_{B}(t)I_{B}(t)\end{aligned}$ *
$\begin{aligned} y_{1}(t)&=K_{A}\widetilde{I}_{A}(t)\\ y_{2}(t)&=K_{B}\widetilde{I}_{B}(t)\end{aligned}$ $\beta_{A},\beta_{B},k,K_{A},K_{B},\tau_{R},\tau_{I},\tau_{P}$ are all locally identifiable.
Behaviorally-coupled diseases Imperfect spillover ($0<s<1$) $\begin{aligned} y_{1}(t)&=I_{A}(t)\\ y_{2}(t)&=I_{B}(t)\end{aligned}$ $\beta_{A},\beta_{B},k,\tau_{R},\tau_{I},\tau_{P}$ are all locally identifiable.
$\begin{aligned} y_{1}(t)&=\beta_{A}S_{A}(t)I_{A}(t)\\ y_{2}(t)&=\beta_{B}S_{B}(t)I_{B}(t)\end{aligned}$ *
$\begin{aligned} y_{1}(t)&=K_{A}\widetilde{I}_{A}(t)\\ y_{2}(t)&=K_{B}\widetilde{I}_{B}(t)\end{aligned}$ $\beta_{A},\beta_{B},k,K_{A},K_{B},\tau_{R},\tau_{I},\tau_{P}$ are all locally identifiable.
The output of STRIKE-GOLDD 4.2 indicates that all the parameters are locally identifiable for prevalence and recognized prevalence. All parameters are locally identifiable for incidence when there is no spillover. We note that when incidence was used as an output measure in the cases for perfect spillover and imperfect spillover, MATLAB ran out of memory and did not produce any results. Thus, we do use incidence as an output measure when performing practical identifiability. All of the initial conditions of the state variables are identifiable except in the case of perfect spillover when using prevalence as the output measure.
3.3.2 Practical Identifiability
When fitting to data, it is important to verify if the model is formulated in a way that allows for the model parameters to be uniquely identified in the context of imperfect data to determine practical identifiability. Using prevalence data and recognized prevalence data, we considered ten scenarios to determine which parameters are strongly identifiable, weakly identifiable, or nonidentifiable. We say that a parameter is strongly identifiable if the average relative error (ARE) for all noise levels satisfies 0 $\leq$ ARE $\leq\sigma$, where $\sigma$ is the measurement error added to the data set. A parameter is weakly identifiable if $\sigma<$ ARE $\leq 10\times\sigma$ for all noise levels, and a parameter is practically nonidentifiable if ARE $10\times\sigma$ for at least one noise level. The ARE values were generated through the use of the Monte Carlo (MC) approach, a sampling method that uses probability distributions to test the practical identifiability of a model (outlined in A1.3) [61, 45]. We use the parameter values in Table 2 and set the parameters $K_{A}$ and $K_{B}$ both equal to 0.5.
The MC approach indicates that for prevalence as an output measures, regardless of the level of spillover, infectivity rates ($\beta_{A}$, $\beta_{B}$) and the delay to adjust risk perception ($\tau_{P}$) are identifiable while sensitivity to risk ($k$) and infection period ($\tau_{I}$) are weakly identifiable. For both $k$ and $\tau_{I}$, the MC approach indicates the parameters are close to being identifiable, as ARE exceeds the noise level ($\sigma$) for only a few instances (see Tables A1, A2, A3, A4, A5). The immunity period ($\tau_{R}$) ranges from weakly identifiable to non-identifiable depending on the level of spillover. For all the scenarios, $\tau_{R}$ is the only parameter in which we observe a drastic jump in the ARE value. For example, there is a 184.73% increase of the ARE value for $\tau_{R}$ when moving from 5% to 10% noise level in Table 7. As we only considered data across a single year and the immunity period is assumed to be 100 days, it is possible that identifiability results would improve for longer periods of data.
For recognized prevalence as an output measure, regardless of the level of spillover, the infectivity of disease AA ($\beta_{A}$) and the delay to adjust risk perception ($\tau_{P}$) are identifiable. Additionally, the infectivity of disease $B$ ($\beta_{B}$), the sensitivity to risk ($k$), infection period ($\tau_{I}$), and the fraction of reports cases for both diseases ($K_{A}$, $K_{B}$) are weakly identifiable across all spillover values tested. For the weakly identifiable parameters, the MC approach indicates the parameters are close to being identifiable, as ARE exceeds the noise level ($\sigma$) for only a few instances (see Tables A6, A7, A8, A9, A10). For recognized prevalence, the immunity period ($\tau_{R}$) is only weakly identifiable for no spillover ($s=0$) and non-identifiable for any amount of spillover. Similar to the prevalence scenario, $\tau_{R}$ is the only parameter that exhibits a substantial increase in the ARE value as the noise level escalates. In Table A6, the ARE value for $\tau_{R}$ increases by 361.92% from noise level 5% to 10%.
Table 7: Practical identifiability for System LABEL:eqn:SIRS with prevalence data under various spillover values. Results obtained using MC approach. See Tables A1, A2, A3, A4, A5 for average relative error (ARE) values used to determine practical identifiability.
Spillover $\beta_{A}$ $\beta_{B}$ $k$ $\tau_{R}$ $\tau_{I}$ $\tau_{P}$
$s=0$ Yes Yes Weakly Weakly Weakly Yes
$s=0.1$ Yes Yes Weakly Weakly Weakly Yes
$s=0.5$ Yes Yes Weakly No Weakly Yes
$s=0.9$ Yes Yes Weakly No Weakly Yes
$s=1$ Yes Yes Weakly Weakly Weakly Yes
Table 8: Practical identifiability for System LABEL:eqn:SIRS with recognized prevalence data under various spillover levels. Results obtained using MC approach. See Tables A6, A7, A8, A9, A10 for average relative error (ARE) values used to determine practical identifiability.
Spillover $\beta_{A}$ $\beta_{B}$ $k$ $\tau_{R}$ $\tau_{I}$ $\tau_{P}$ $K_{A}$ $K_{B}$
$s=0$ Yes Weakly Weakly Weakly Weakly Yes Weakly Weakly
$s=0.1$ Yes Weakly Weakly No Weakly Yes Weakly Weakly
$s=0.5$ Yes Weakly Weakly No Weakly Yes Weakly Weakly
$s=0.9$ Yes Weakly Weakly No Weakly Yes Weakly Weakly
$s=1$ Yes Weakly Weakly No Weakly Yes Weakly Weakly
4 Discussion
This paper examines the dynamics of interdependencies across multiple pathogens in the absence of cross-immunity. While previous studies have often attributed inter-pathogen interactions to cross-immunity – where infection by one pathogen confers immunity against another [6, 24] – this study relaxes that assumption and instead highlights a behavioral mechanism that can couple the spread of two pathogens. This perspective is motivated by observations during the COVID-19 pandemic, where a surge in COVID-19 cases coincided with a marked decline in Influenza A and B cases [59, 30]. This decline was largely due to public adherence to NPIs implemented to control COVID-19, which also inadvertently suppressed the spread of influenza [62]. This observation raises an important question: under what conditions can behavioral responses to one deadly disease contain the spread of others? We do not aim to reproduce the full complexity of epidemic dynamics here; rather, this work is theoretical investigation into behaviorally mediated interactions between diseases in a framework where analysis is possible.
To this end, we develop a mathematical model to examine behavioral interdependencies across two pathogens. Building on a previously established behavioral epidemic model (SEIRb) [37, 57], we extend the framework to capture the spread of two pathogens, where behavioral responses to one can influence the transmission dynamics of the other. We refer to this form of interdependence as behavioral spillover. Our model represents two diseases of different infectivity, and varying degrees of behavioral interdependence. Specifically, the model investigates the extent to which NPIs targeting one pathogen affect the spread of another.
Our analysis explores a range of simulation outcomes for different values of the basic reproduction number and spillover strength. The results indicate that NPIs targeted at a pathogen with a relatively high basic reproduction number can significantly suppress the spread of other pathogens in the short term. This leads to scenarios in which one pathogen exceeds the rest while others maintain effective reproduction numbers below one, i.e., are fully suppressed. For example, in the presence of complete behavioral spillover, we observed that the population may only experience the disease with the highest $\mathcal{R}_{0}$, as if the ecosystem, including the diseases and humans, selects the pathogen with the highest $\mathcal{R}_{0}$. Additionally, we identify parameter regimes in which both pathogens coexist but spread at reduced rates compared to their behavior in isolation. With imperfect spillover, where NPIs designed for one disease only partially affect the spread of others, we estimate the range of values $\mathcal{R}_{0}$ in which only one of the pathogens spreads.
Moreover, our study explored equilibria properties of the new model given in System (LABEL:eqn:SIRS). Each scenario has four possible equilibria, depending on the values of the $\mathcal{R}_{0,i}$: one disease-free, two boundary (endemic in one disease and disease-free in the other), and one endemic equilibrium. While results regarding the disease-free equilibria were typical for disease models, i.e., the basic reproductive number functioned as a threshold value at one, results for the boundary and endemic equilibria when behavioral spillover was present diverged from this norm as one was not the threshold for stability for the equilibria. Within these scenarios, it is possible to have $\mathcal{R}_{0,A},\mathcal{R}_{0,B}>1$ with a stable boundary equilibrium and unstable endemic equilibrium; that is, the disease with the greater reproduction number will persist, while the other disease will die out despite both having a basic reproduction number above one, the typical threshold for persistence [7, 36, 26]. In cases with spillover, transmission depends on both diseases; analytically, the threshold condition for the newly introduced disease thus depends on the value of the infectious equilibrium of the endemic disease. If this threshold is large enough (i.e. if the presence of the endemic disease is large enough), then the newly introduced disease will be unsuccessful in establishing itself endemically. We identify the thresholds that determine the persistence disease(s) for all three scenarios.
As a critical component of working with a novel model is the parameterization with respect to real world data. This procedure usually involves a thorough literature review or model fitting. To assess the potential utility of our novel model to be fit to data, we conducted a structural identifiability analysis [61, 67]. Our identifiability analysis demonstrated that all the parameters are structurally locally identifiable with respect to prevalence and recognized prevalence data. The only scenario in which we obtained unidentifiable results was for the initial conditions of perceived prevalence, $\widetilde{I}_{A}(0)$ and $\widetilde{I}_{B}(0)$, when using prevalence data in the case of perfect spillover. Although $\widetilde{I}_{A}(0)$ and $\widetilde{I}_{B}(0)$ are unidentifiable, the initial condition values are known, which provided confidence to proceed with the practical identifiability analysis. From the practical identifiability analysis, we observed that the majority of the parameters were identifiable in some form (i.e. strongly or weakly) with respect to prevalence and recognized prevalence. The immunity period, $\tau_{R}$, was the only parameter that was not identifiable. Since we used $\tau_{R}=100$ representing an immunity period of 100 days, and we simulated the diseases dynamics for 100 days, it is probable that $\tau_{R}$ is identifiable if the simulations were lengthened considerably or if the immunity period was shortened.
This study contributes to the fields of mathematical biology and epidemiological modeling, particularly to the emerging area of behavioral epidemic modeling [21]. Firstly, it adds to the body of literature on multi-pathogen dynamics [6, 24] by offering a contrasting perspective: that interdependent epidemic patterns can arise even in the absence of cross-immunity, purely due to behavioral interdependencies. This paper introduced the term behavioral spillover to describe this phenomenon, in which behavioral responses to one pathogen influence the transmission of others. An important implication is that ignoring such behavioral interdependencies may lead to misleading conclusions—such as over- or under-estimating pathogen transmission rate or wrongly inferring the presence of cross-immunity between pathogens. Second, this study contributes to the endogenous formulation of human behavior [38]. It builds on prior behavioral epidemic models that incorporate a risk-response feedback loop, wherein rising (or falling) prevalence increases (or decreases) adherence to non-pharmaceutical interventions (NPIs), which in turn influences transmission [56, 20, 42]. The paper extends this framework to demonstrate how multiple pathogens can become dynamically coupled through shared behavioral responses.
However, this study has several limitations that point to directions for future research. First, the current model assumes a homogeneous population and does not account for heterogeneity in risk or behavioral response. For example, while elderly populations are at higher risk for both COVID-19 and Influenza A/B, other diseases like Lyme disease primarily affect different demographics, such as children. Incorporating population heterogeneity could yield more nuanced insights. Second, the model does not include vaccination dynamics, which can significantly alter behavioral interdependencies. For instance, after COVID-19 vaccines became available, increased perceived safety led to greater social activity, potentially raising the risk of influenza spread. Accounting for such vaccine-induced behavioral shifts would be a valuable extension. Finally, the model assumes infected individuals do not alter their behavior and does not include biological cross-immunity between pathogens. While these assumptions simplify the system, they allow us to isolate behavioral mechanisms, showing that apparent protection against one disease can arise purely from changes in perceived risk of the other. Relaxing these assumptions—by incorporating quarantine upon infection or partial cross-immunity—could further strengthen the interdependencies and accelerate the practical extinction of diseases with lower reproduction numbers.
In conclusion, while the current model focuses on two pathogens for clarity, it can be generalized to consider multiple specific pathogens, different variants, and combinations of cross-immunity and behavioral spillover. Such extensions could offer broader insights into epidemic interactions in complex real-world settings. Furthermore, our focus was on offering a mathematical formulation of relevant models and the effort should be extended to include empirical tests using data across different regions and over a long period of time.
5 Disclosure Statement
The authors report there are no competing interests to declare.
6 Funding
This work was supported by the US National Science Foundation, Division of Mathematical Sciences and Division of Social and Economic Sciences under Grant # 2229819.
7 Data Availability Statement
All code files will be posted to GitHub upon acceptance. A zipped file containing code is available for review.